OpenStudy (anonymous):

find the x and y intercepts of f. f(x)= (x-2)^2 * (x^2 - 25)

OpenStudy (anonymous):

Please help walk me through it, I am reviewing and don't remember how to do this at all

OpenStudy (anonymous):

for y intercepts, set x=0 and calculate value of f(x), for x intercepts, set f(x)=0 and calculate values of the solution for f(x)=0

OpenStudy (anonymous):

I know that one y intercept is 2. Do the nonreal intercepts from the (x^2 - 25) count? Do I need to find those?

OpenStudy (anonymous):

non real intercepts do not count but where are you getting non real solutions for f(x)=0?

OpenStudy (anonymous):

(x-2)^2(x^2-25)=(x-2)(x-2)(x+5)(x-5)=0

OpenStudy (anonymous):

duh sorry I got myself confused... :O

OpenStudy (anonymous):

you got it right?

OpenStudy (anonymous):

I got the y intercept but the x intercept never did make sense to me... um how do I do that?

OpenStudy (anonymous):

oh wait do you just plug in 0 for x everywhere?

OpenStudy (anonymous):

y intercept is the point where where the graph cuts y axis, you know this right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

on the y axis, you must know x=0, right?

OpenStudy (anonymous):

so you just plug in 0 for x? I learned this last fall so it is slowly coming back...

OpenStudy (anonymous):

yeah, correct

OpenStudy (anonymous):

so basically you are calculating the value of f(x) at x=0 as you are plotting f(x) at y direction, value of f(x) at x=0 is the y intercept

OpenStudy (anonymous):

thank you I just needed that refresher

OpenStudy (anonymous):

you are welcome :)

OpenStudy (anonymous):

Have a great day!